OpenStudy (anonymous):

Which shows three ordered pairs that are all solutions of the equation y = x + 7? A. (0, 7), (2, 9), and (3, 11) B. (0, 7), (2, 9), and (3, 10) C. (1, 7), (2, 9), and (3, 10) D. (1, 8), (2, 9), and (3, 11)

OpenStudy (sleepyjess):

Just substitute in the options for x and y :)

OpenStudy (acxbox22):

plug all the coordinates for each option into the equation and see if the equation holds true

OpenStudy (anonymous):

what

OpenStudy (sleepyjess):

Let me explain. In the first coordinate of the first option, 0 is x and 7 is y. So, if you substitute those in, you get 7 = 0 + 7

OpenStudy (acxbox22):

ex. A) (0,7) >>>> 7=0+7 >>> 7=7 << correct now do that for all the coordinates

OpenStudy (acxbox22):

@sleepyjess keeps on beating me.....

OpenStudy (anonymous):

so its a

OpenStudy (sleepyjess):

Which is true, now can you try for the coordinate 2, 9?

OpenStudy (acxbox22):

no i am just giving an example

OpenStudy (sleepyjess):

No, both A and B have the coordinate 0, 7. So we need to test the others.

OpenStudy (anonymous):

still dont get it

OpenStudy (sleepyjess):

hmmmm... what don't you understand?

OpenStudy (anonymous):

everything

OpenStudy (sleepyjess):

ok, do you understand what a coordinate is?

OpenStudy (anonymous):

a little

OpenStudy (anonymous):

i think its b

OpenStudy (sleepyjess):

A coordinate is an x and y value, in the order x, y. For example, in two of the options, there is the coordinate 0, 7. 0 is x, and 7 is y.

OpenStudy (anonymous):

oh ok

OpenStudy (sleepyjess):

Do you understand that so far?

OpenStudy (anonymous):

yea but i need the answer and how u got it bc i am takeing a test

OpenStudy (sleepyjess):

If this is a test, then I can't help. Sorry :/

OpenStudy (anonymous):

ugh

OpenStudy (sleepyjess):

Asking and answering test questions is against OpenStudy's Anti-Cheating Policy and your school's honor code.

OpenStudy (anonymous):

ugh why did i even waist my time